# Modelling Sampling Distributions of Test Statistics with Autograd

### v1

@@ 4.2 Uncertainty quantification using Bayesian neural networks
In practice, given the computational burden of sampling from the true posterior density, we adopt a simpler approach. Let $\mathbf{w^{*}}\in\mathbb{R}^{k}$ be the best-fit neural network parameters obtained by our training protocol. We fluctuate the neural network parameters by sampling $\mathbf{w}_{1},...,\mathbf{w}_{n}\sim\mathcal{N}(\mathbf{w^{*}},\sigma^{2}\mathbf{I}_{k})$ for different choices of $\sigma$. With $n=100$, the resulting ensemble of models evaluated at the sampled parameters $\hat{f}(\mathbf{w}_{1}),...,\hat{f}(\mathbf{w}_{n})$ are shown in Fig. 7.
4.3 Bootstrap NN
We follow the bootstrap idea (see Sec. 2.5) to draw datasets with replacement from the training data, such that each bootstrap sample has the same size as the original training set. This is done $K=200$ times, resulting in $K=200$ bootstrap training datasets. The same model with the architecture described in Sec. 4.1 is re-trained for $10^{5}$ iterations, with each model fitted to a different bootstrap sample. The spread of the resulting models over the $K=200$ neural networks is displayed in Fig. 8. The envelope of the spread is observed to be much smaller than for the previous methods. This could be due to insufficient training of the models. Other possible reasons for this are outlined below.
@@ 4.3 Bootstrap NN
Although the bootstrap has been hailed as a satisfactory method for uncertainty quantification due to its theoretical guarantees [47] it has limitations. For example, it is known to fail in situations where the data have a non-trivial dependency structure or when the functionals of interest are not smooth [48].
@@
Another question that was recently explored is the efficacy of bootstrap in the high-dimensional regime $n/d<1$ relevant to machine learning, where $d$ is the dimension of the machine learning model parameter space and $n$ the training sample size. In [49] it is shown that even for the simple case of inference in the linear model with unregularized least squares, bootstrap techniques perform poorly in only moderately high dimensions even with $n>d$. It was also shown in [50] that resampling methods such as the bootstrap yield reliable error estimates only in the very low-dimensional regime $n>>d$, and is fraught with problems in high-dimensional regimes. Therefore, perhaps one shouldn’t be surprised by the results described above.
5 Discussion and conclusion
@@
In this paper we try to answer the question whether one can accurately model the sampling distribution of a test statistic starting from a neural network model of the cdf.  [^3] Code to reproduce all of our results is available at https://github.com/hbprosper/cdf2pdf.. Our motivation for modelling the cdf $F(\lambda\mid\boldsymbol{\theta})$ as opposed to an ab initio modelling of $f(\lambda\mid\boldsymbol{\theta})$ is that in general a better precision is reached approximating integrals of functions rather than the functions themselves. Furthermore, since the step of taking the derivative with autograd is exact, the accuracy of the approximated cdf directly translates to the accuracy of the desired pdf.
@@
It was found that a recently introduced method (ALFFI) that approximates the cdf as the mean of a certain discrete random variable appears not to be accurate enough in modelling the cdf for the purpose of deriving its associated pdf. We caution however that this may be the inevitable consequence of trying to model the cdf of an intrinsically discrete distribution with a smooth approximation.
@@
We employ relatively simple yet effective neural network models such that the amount of fine-tuning that is involved in going to more complex data settings is minimal. Several neural network uncertainty quantification techniques were reviewed and implemented and we studied possible corrections of the cdf. Of the uncertainty quantification methods considered only one is calibrated by construction, namely, conformal inference. We recommend that it be used as a simple benchmark with respect to which other uncertainty quantification methods can be compared and calibrated.
@@
As can be seen in Eq. 4 conformal inference results in confidence sets that guarantee marginal coverage, $\mathbb{P}\left\{y_{n+1}\in\mathcal{C}_{\alpha}\left(x_{n+1}\right)\right\}\geq 1-\alpha$, that is, the coverage marginalized (integrated) over the population of which a dataset is a sample. That property is not at all controversial as coverage is a property of the population from which the dataset is presumed to have been sampled. But if each element $x$ of the dataset contains a subset that are parameters $\theta$ sampled from a prior $\pi_{\theta}$ then the marginalization is also with respect to this prior. When a prior is involved this form of coverage is weaker than the more sought after conditional coverage guarantee $\mathbb{P}\left\{y_{n+1}\in\mathcal{C}_{\alpha}\left(x_{n+1}\right)\mid\theta\right\}\geq 1-\alpha$. It was shown in [23] that such distribution-free conditional coverage is impossible to achieve with a finite sample, which has inspired studies that try to bridge the gap between marginal and conditional coverage [51]. In our case, $x=\{\theta,\lambda\}$ and we generate calibration sets at each $\theta$ point, thereby making our conformal confidence sets adaptive, that is, change with $\theta$. However, they are still not adaptive with respect to $\lambda$; the conformal confidence intervals are the same width at each $\lambda$ value.
@@
Appendix A ALFFI algorithm for modelling CDF for ON/OFF Example
ALFFI algorithm for modelling CDF in the On/Off problem, inspired by [7].
@@ Appendix A ALFFI algorithm for modelling CDF for ON/OFF Example
16:Use $\mathcal{T^{\prime}}$ to learn the parameterized function ^$\widehat{\mathbb{C}}(\lambda\mid\mu,\nu):=\mathbb{E}[Z\mid\mu,\nu]$ via regression of $Z_{i}$ on $\{\mu_{i},\nu_{i},\lambda_{i}\}$ using mean square error as the loss function
@@ Appendix B Residuals
Figure 9: $\Delta C=\hat{q}-p$ residuals as a function of $\hat{q}$ at different $(\mu,\nu)$ points.
Figure 10: $\Delta r=\hat{F}-F$ a a function of $\lambda$ at different $(\mu,\nu)$ points.

### v2

@@ 4.2 Uncertainty Quantification using Bayesian Neural Networks
In practice, given the computational burden of sampling from the true posterior density, we adopt a simpler approach. Let $\mathbf{w^{*}}\in\mathbb{R}^{k}$ be the best-fit neural network parameters obtained by our training protocol using the Huber loss. We fluctuate the neural network parameters by sampling $\mathbf{w}_{1},...,\mathbf{w}_{n}\sim\mathcal{N}(\mathbf{w^{*}},\sigma^{2}\mathbf{I}_{k})$ for different choices of $\sigma$. With $n=100$, the resulting ensemble of models evaluated at the sampled parameters $\hat{f}(\mathbf{w}_{1}),...,\hat{f}(\mathbf{w}_{n})$ are shown in Fig. 12.
4.3 Bootstrap Neural Networks
We follow the bootstrap idea (see Sec. 2.6) to draw datasets with replacement from the training data, such that each bootstrap sample has the same size as the original training set. This is done $K=200$ times, resulting in $K=200$ bootstrap training datasets. Due to practical considerations, we adopted a simplified neural network architecture consisting of 5 hidden layers, each containing 10 neurons, followed by a single output neuron utilizing a sigmoid activation function. The SiLU activation function was employed at each hidden node to enhance nonlinearity and learning capabilities. The network was trained using the NAdam optimizer with a fixed learning rate of $3\times 10^{-4}$ in mini batches of size 60. This model architecture was retrained for $10^{5}$ iterations, with each model fitted to a different bootstrap sample.
The spread of the resulting models over the $K=200$ neural networks is displayed in Fig. 13. The envelope of the spread is observed to be much smaller than for the previous methods. This could be due to insufficient training of the models. Other possible reasons for this are outlined below.
@@ 4.3 Bootstrap Neural Networks
Although the bootstrap has been hailed as a satisfactory method for uncertainty quantification due to its theoretical guarantees [50] it has limitations. For example, it is known to fail in situations where the data have a non-trivial dependency structure or when the functionals of interest are not smooth [51]. Another question that was recently explored is the efficacy of bootstrap in the high-dimensional regime $n/d<1$ relevant to machine learning, where $d$ is the dimension of the machine learning model parameter space and $n$ the training sample size. In [52] it is shown that even for the simple case of inference in the linear model with unregularized least squares, bootstrap techniques perform poorly in only moderately high dimensions even with $n>d$. It was also shown in [53] that resampling methods such as the bootstrap yield reliable error estimates only in the very low-dimensional regime $n\gg d$, and is fraught with problems in high-dimensional regimes. Therefore, perhaps one shouldn’t be surprised by the results described above.
5 Discussion and Conclusions
@@
Automatic differentiation is a fundamental tool in training neural networks, efficiently computing gradients with respect to model parameters and thus facilitating optimization algorithms like gradient descent. In this work, we exploit this capability to compute derivatives with respect to input variables just as is done, for example, in physics-informed neural networks (see for example [54] and references therein). This allows one to investigate whether it is possible to accurately model a pdf by starting from a neural network model of the associated cdf. Cumulative distribution functions naturally arise in frequentist simulation-based inference. One motivation for modeling the cdf $F(\lambda\mid\boldsymbol{\theta})$ as opposed to an ab initio modeling of $f(\lambda\mid\boldsymbol{\theta})$ is that in general a better precision is reached approximating integrals of functions rather than the functions themselves. Furthermore, since the step of taking the derivative with autograd is exact, the accuracy of the approximated cdf directly translates to the accuracy of the desired pdf.
@@
It was found that a recently introduced method (ALFFI) that approximates the cdf as the mean of a certain discrete random variable appears not to be accurate enough in modeling the cdf for the purpose of deriving its associated pdf. We caution however that this may be the inevitable consequence of trying to model the cdf of an intrinsically discrete distribution with a smooth approximation. Our choice to investigate this distribution is intentional: it is a difficult test case for the proposed method.
@@
We employ relatively simple yet effective NN models such that the amount of fine-tuning that is involved in going to more complex data settings is minimal. Several NN uncertainty quantification techniques were reviewed and implemented and we studied possible corrections of the cdf. Of the uncertainty quantification methods considered only one is calibrated by construction, namely, conformal inference. We recommend that it be used as a simple benchmark with respect to which other uncertainty quantification methods can be compared and calibrated.
@@
As can be seen in Eq. 5 conformal inference results in confidence sets that guarantee marginal coverage, $\mathbb{P}\left\{y_{n+1}\in\mathcal{C}_{\alpha}\left(x_{n+1}\right)\right\}\geq 1-\alpha$, that is, the coverage marginalized (integrated) over the population of which a dataset is a sample. That property is not at all controversial as coverage is a property of the population from which the dataset is presumed to have been drawn. But if each element $x$ of the dataset contains a subset that are parameters $\theta$ sampled from a prior $\pi_{\theta}$ then the marginalization is also with respect to this prior. When a prior is involved this form of coverage is weaker than the more sought after conditional coverage guarantee $\mathbb{P}\left\{y_{n+1}\in\mathcal{C}_{\alpha}\left(x_{n+1}\right)\mid\theta\right\}\geq 1-\alpha$. It was shown in [24] that such distribution-free conditional coverage is impossible to achieve with a finite sample, which has inspired studies that try to bridge the gap between marginal and conditional coverage [55]. In our case, $x=\{\theta,\lambda\}$ and we generate calibration sets at each $\theta$ point, thereby making our conformal confidence sets adaptive, that is, change with $\theta$. However, they are still not adaptive with respect to $\lambda$; the conformal confidence intervals are the same width at each $\lambda$ value. But it should be noted that this is not necessarily unusual. Consider, for example, the 68.3% confidence interval $[x-\sigma,x+\sigma]$ of a Gaussian pdf. The widths of these intervals are twice the standard deviation $\sigma$ and, therefore, independent of $x$.
@@
Data Availability All data and code to reproduce all of our results is available at https://github.com/hbprosper/cdf2pdf.
[…]
@@
Appendix A ALFFI algorithm for modeling CDF for ON/OFF Example
ALFFI algorithm for modeling CDF in the ON/OFF problem, inspired by [7].
@@ Appendix A ALFFI algorithm for modeling CDF for ON/OFF Example
16:Use $\mathcal{T^{\prime}}$ to learn the parameterized function ^$\widehat{\mathbb{C}}(\lambda\mid\mu,\nu):=\mathbb{E}[Z\mid\mu,\nu]$ via regression of $Z$ on $\{\mu,\nu,\lambda\}$ using mean square error as the loss function
@@ Appendix B Residuals
Figure 14: “quantile residuals” $\Delta C=\hat{q}-p$ as a function of $\hat{q}$ at different $(\mu,\nu)$ points.
Figure 15: Cdf residuals $\Delta r=\hat{F}-F$ a a function of $\lambda$ at different $(\mu,\nu)$ points.
Declarations
[…]
Conflict of Interest All authors declare that they have no conflict of interest.